OpenStudy (anonymous):

6/10% x 45 explain/check

OpenStudy (anonymous):

@DanJS

OpenStudy (anonymous):

hold on let me solve it first

OpenStudy (danjs):

6 tenths of a percent or (6/10) percent

OpenStudy (danjs):

as in 60 percent

OpenStudy (anonymous):

so if the denominator is 10 do i still multiply by 100?

OpenStudy (danjs):

Is it six tenths of a percent as in 0.6/100 or (6/10) percent as in 60/100?

OpenStudy (anonymous):

idk im lost now

OpenStudy (danjs):

assuming it is \[\frac{ 6 }{ 10 }\]%

OpenStudy (danjs):

6 out of 10 percent of 54

OpenStudy (danjs):

or 60% of 45

OpenStudy (anonymous):

it just says 6/10 %

OpenStudy (danjs):

\[\frac{ 6*10 }{ 10*10 } * 45 = \frac{ 60 }{ 100 }*45 =\]

OpenStudy (anonymous):

27??

OpenStudy (anonymous):

.027 i meant

OpenStudy (danjs):

Reason it out, what number is 60% of 45? 27 or .027?

OpenStudy (danjs):

If you got a 60% on a test worth 4 points, did you get less than 1 point or 27 points

OpenStudy (anonymous):

27 but the answer choices dont have just 27

OpenStudy (danjs):

45 points*

OpenStudy (danjs):

hmm, then it must be just 6/10 of a percent , or \[\frac{ 0.6 }{ 100 }*45\]

OpenStudy (anonymous):

i really dont know lol

OpenStudy (danjs):

Is one of the answers0.27

OpenStudy (danjs):

6/10ths of a percent of 45 is 0.27